# Scalable LinUCB: Low-Rank Design Matrix Updates for Recommenders with Large Action Spaces

### v1

@@ Abstract
Linear contextual bandits, especially LinUCB, are widely used in recommender systems. However, its training, inference, and memory costs grow with feature dimensionality and the size of the action space. The key bottleneck becomes the need to update, invert and store a design matrix that absorbs contextual information from interaction history. In this paper, we introduce Scalable LinUCB, the algorithm that enables fast and memory efficient operations with the inverse regularized design matrix. We achieve this through a dynamical low-rank parametrization of its inverse Cholesky-style factors. We derive numerically stable rank‑1 and batched updates that maintain the inverse without directly forming the entire matrix. To control memory growth, we employ a projector‑splitting integrator for dynamical low‑rank approximation, yielding average per‑step update cost $\mathcal{O}(dr)$ and memory $\mathcal{O}(dr)$ for approximation rank $r$. Inference complexity of the suggested algorithm is $\mathcal{O}(dr)$ per action evaluation. Experiments on recommender system datasets demonstrate the effectiveness of our algorithm.
@@
1 Introduction
Interactive recommender systems must learn from sparse and delayed feedback while responding within tight latency budgets. A popular class of methods for solving this problem are the contextual bandits, which are gaining increasing popularity due to their ability to fine-tune on new data and efficiently manage the exploration–exploitation trade-off auer2002finite . These algorithms are not only well-studied theoretically, with strong finite-time guarantees auer2002finite , abbasi2011improved , agrawal2013thompson , but are also widely used in production recommender systems li2010contextual , yi2023online .
Within this class of methods, one of the most popular algorithms is the well-known LinUCB li2010contextual . The model is simple, its parameters are interpretable, and the update rule is easy to implement. LinUCB assumes a linear relationship between the $d$-dimensional context vector and the expected reward of the selected action. Regret guarantees for the LinUCB algorithm are well studied abbasi2011improved ; moreover, LinUCB admits practical implementations for online recommendations li2010contextual . These properties make it a reasonable baseline when a system requires online updates.
Scalability becomes an issue as the feature dimension $d$ and the number of items (actions) grow. Exact implementations Angioli2025Efficient maintain per-arm design matrices with ridge regularization and apply rank-1 updates using the Sherman–Morrison identity. This keeps training at $\mathcal{O}(d^{2})$ per round but still requires $\mathcal{O}(d^{2})$ memory per arm and inversion of the design matrix for each candidate to evaluate confidence widths (exploration bonuses). With large action sets, inference becomes prohibitively costly since all items (actions) must be scored per request. Prior works reduce these costs by sketching the design matrix (e.g., Frequent Directions and its compensated/adaptive variants) kuzborskij2019efficient , chen2020efficient , wen2024matrix , by projecting features to a lower dimension yu2017cbrap , or by using diagonal/block-diagonal approximations yi2023online . Each option involves a different trade-off between accuracy, robustness, and computational complexity.
In this paper, we propose a scalable version of the LinUCB algorithm based on dynamical updates of Cholesky-style symmetric factors of the inverse regularized design matrix. Our algorithm enables fast training and inference while preserving the quality of the original LinUCB algorithm. At the same time, it reduces the memory consumption and iteration cost of the method. Our primary contributions are as follows:
  • We maintain and dynamically update a Cholesky-style factor representation for the inverse regularized design matrix in the LinUCB algorithm. We derive the corresponding dynamical low-rank update, which avoids explicit computation of inverse matrices. We employ a projector-splitting method lubich2014projector for dynamical low-rank approximation to cap the approximation rank and obtain predictable time and memory costs, referring to the proposed method as PSI-LinUCB. The proposed algorithm yields an average per-interaction update cost of $\mathcal{O}(dr)$ during training, where $r$ is the approximation rank. The inference complexity of the PSI-LinUCB algorithm is $\mathcal{O}(dr)$ per action evaluation.
  • We evaluate our method on the classical MovieLens-1M dataset as well as on a set of large-scale recommendation datasets from the Amazon collection — the setting in which standard LinUCB immediately fails even on top-grade hardware due to scalability issues. The proposed method achieves a $8.8\times$ reduction in training and inference time with respect to the standard LinUCB algorithm with Sherman–Morrison updates Angioli2025Efficient . It also reduces memory consumption in $15.9\times$ while maintaining the quality metric (hit rate) at the level of the vanilla LinUCB implementation.
2 Problem setting
We begin this section with describing the LinUCB algorithm li2010contextual , arguably the most popular algorithm among the class of contextual bandits. This algorithm employs the Upper Confidence Bound (UCB) exploration strategy, following the principle of optimism in face of uncertainty auer2002finite . Theoretical properties of LinUCB algorithm attracted a vast amount of contributions abbasi2011improved , chu2011contextual , valko2013finite , lattimore2020bandit .
@@ 2 Problem setting
Formally, at each time step $t\in\{0,\ldots,T\}$, agent observes a set of arms (actions) $\mathcal{A}_{t}$, where each arm $a\in\mathcal{A}$ is associated with a context vector ${x}_{t,a}\in\mathbb{R}^{d}$. LinUCB assumes a linear relationship between the context vector and the reward, that is, expected reward for arm $a$ at time $t$ is given by
@@
$a_{t}=\arg\max_{a\in\mathcal{A}_{t}}\left(\hat{{\theta}}_{a}^{\top}{x}_{t,a}+\alpha\sqrt{{x}_{t,a}^{\top}A_{a}^{-1}{x}_{t,a}}\right),$ (6)
@@
where $\alpha\in\mathbb{R}$ is a confidence parameter controlling the exploration–exploitation trade-off. In online learning problems $\alpha>0$, but sometimes learning problems with fixed dataset require setting negative values of $\alpha$ (the so-called anti-exploration), see e.g. rezaeifar2022offline , xu2024provably .
In the context of recommender systems, at each time step $t$, the agent observes a particular user $u$ and a set of available items $\mathcal{A}_{t}$. For consistency with the previous notation, the context vector $\mathbf{x}_{t,a}$ can be equivalently written as $\mathbf{x}_{u,a}$, representing the joint feature vector of the user–item pair $(u,a)$, which captures both user-specific and item-specific characteristics.
The LinUCB algorithm described above can be generalized to the setting with batch updates, where model parameters are updated after multiple interactions (after a batch of interactions) rather than once over the whole training dataset, see Algorithm˜1.
Algorithm 1 LinUCB
1: for each arm $a\in\mathcal{A}$ set $A_{0,a}=\lambda I$
2:for t in {0,…,n-1} do
3:  $X_{t,a}=[]$, $R_{t,a}=[]$
4:  for all $(u,a,r)$ in $batch_{t}$ do
5:   $X_{t,a}.append\,[x_{u,a}]$
6:   $R_{t,a}.append\,[r]$
7:  end for
8:  for each arm $a$ in $batch_{t}$ do
9:   $A_{t+1,a}\leftarrow A_{t,a}+X_{t,a}^{\top}X_{t,a}$
10:   $b_{t+1,a}\leftarrow b_{t,a}+X_{t,a}^{\top}R_{t,a}$
11:   ${\theta}_{t+1,a}\leftarrow A_{t+1,a}^{-1}b_{t+1,a}$
12:  end for
13:end for
14:for each arm $a\in\mathcal{A}$ do
15:  Observe context $x_{u,a}$
16:  Compute $\text{UCB}_{u,a}={\theta}_{n,a}^{\top}{x}_{u,a}+\alpha\sqrt{{x}_{u,a}^{\top}{A}_{n,a}^{-1}{x}_{u,a}}$
17:end for
18:return $\arg\max_{a\in\mathcal{A}}\text{UCB}_{u,a}$
However, the LinUCB algorithm has several scalability challenges:
@@
  • Matrix inversion: The inversion of the matrix $A_{a}$ has a computational complexity of $O(d^{3})$, which makes the algorithm computationally expensive in high-dimensional context spaces.
@@
  • Large action space: A separate matrix $A_{a}\in\mathbb{R}^{d\times d}$ must be stored and updated for each arm, which not only increases time and memory requirements but also the number of matrix operations as the action space grows.
3 Related works
To address scalability constraints, several extensions and variants of the LinUCB algorithm have been proposed in the literature.
A widely used approach to accelerate LinUCB is to apply rank-1 updates of the ridge-regularized design matrix $A_{a}$ via the Sherman–Morrison identity  Angioli2025Efficient , Ciucanu2022Implementing , Wang2022Dynamic , Yan2025CoCoB , Ozbay2024Comparative . These updates reduce the per-round update complexity from $\mathcal{O}(d^{3})$ to $\mathcal{O}(d^{2})$ while maintaining exact parameter estimates, but still require storing a full $d\times d$ matrix for each arm.
Another line of research aims to improve the efficiency of LinUCB through matrix sketching techniques. Early work kuzborskij2019efficient demonstrated that LinUCB, can be efficiently implemented using the Frequent Directions (FD) sketching method, reducing the per-round update time from $\mathcal{O}(d^{2})$ to $\mathcal{O}(md)$, where $m$ is the sketch size and $d$ is the feature dimension. However, applying FD to contextual bandits has certain drawbacks. In particular, the FD sketching method violate the positive definite monotonicity design matrices $A_{a}$, which may affect stability and theoretical guarantees. To address this limitation, a Spectral Compensation Frequent Directions (SCFD) method and its adaptation for high-dimensional (CBSCFD) algorithms were proposed chen2020efficient , which preserves positive definiteness while maintaining the same $\mathcal{O}(md)$ computational and memory complexity. More recently, adaptive sketching techniques such as Dyadic Block Sketching wen2024matrix have been introduced to dynamically adjust the sketch size, ensuring a per-round update complexity of $\mathcal{O}(dl)$, where $l$ is the current sketch size. This adaptive strategy prevents excessive spectral loss and avoids linear regret when the spectrum of the design matrix decays slowly.
A detailed discussion of the limitations of the approaches reviewed above is provided in Section˜6.1.
@@ 4 Scalable LinUCB with low-rank updates
The main challenges related to LinUCB updates are its scalability with the number of arms $\mathcal{A}$ and the complexity of updating the parameter $\hat{\theta}_{a}$ in (5). Indeed, after observing each batch of interactions, the learner has to update the parameters that correspond to every arm $a$ appeared in the current batch. This update requires computing $A_{a}^{-1}$, which has the complexity of $\mathcal{O}(d^{3})$ in the naive implementation. Moreover, computing the decision-making rule (6) requires matrix-vector multiplications with the same matrix $A_{a}^{-1}$.
A natural idea to reduce complexity of calculating $A_{a}^{-1}$ is to consider the setting of rank-1 updates (when the context vector arrives one at a time) together with the Sherman-Morrison update equation. This approach was implemented in context of recommender systems in Angioli2025Efficient . Further ways to reduce the model complexity is to use a low-rank approximation of the matrix
@@
which is a part of $A_{a}$ in (3). This approach is used in the frequent directions (FD) algorithm kuzborskij2019efficient and CBSTD algorithm chen2020efficient . In contrast to these works, we employ an alternative approach, which relies on the dynamic representation of the inverse matrix $A_{a}^{-1}$. Towards this aim, we use the Cholesky-style symmetric factorization:
$A_{a,t}=L_{a,t}L_{a,t}^{\top}$ (8)
and note that in this case
$A_{a,t}^{-1}=L_{a,t}^{-\top}L_{a,t}^{-1}\;.$ (9)
The representation (9) is a key formula for our further analysis. We acknowledge the following important statement - we do not compute the inverse factors $L_{a,t}^{-1}$ from the matrix $L_{a,t}$, but use a dynamical updates for $L_{a,t}^{-1}$ based on $L_{a,t-1}^{-1}$. Since the update mechanism is identical for each arm, we omit the index $a$ and write $A_{t}$ and $L_{t}$ (respectively, $L_{t}^{-1}$) instead of $A_{a,t}$ and $L_{a,t}$.
In the next parts of this section, we present our methodology in different setups. First, in Section˜4.1 we consider the case of rank-1 updates and write the dynamics of updates of the inverse root matrix $L_{t}^{-1}$. Then in Section˜4.2 we generalize our expressions for the batch update case. Both representations rely on a recursive definition in the form of matrix decomposition $U_{t}V_{t}^{\top}$. In Section˜5, we show that this decomposition can be dynamically updated in a low-rank format using the Projector-Splitting Integrator (PSI) method of Lubich and Oseledets lubich2014projector . Such updating mechanism restrains an uncontrollable growth of factor matrices $U_{t}$ and $V_{t}$ over time and enables “on-the-fly” adaptation to the stream of contextual data. This approach has been previously considered in the context of modifying the PureSVD model in recommender systems Olaleke2021Dynamic .
@@ 4.1 Rank-1 updates
We start by considering the case of context vectors arriving one-by-one requiring the parameter updates to be performed after every interaction. The update to design matrix $A_{t}$ at time $t+1$ can be written in the form of a rank-1 perturbation
$A_{t+1}=A_{t}^{\top}+x_{t+1}x_{t+1}^{\top}.$ (10)
A straightforward approach to calculating updates to the inverse is to apply the Sherman-Morrison formula and obtain the recursive form through $t$ as done by Angioli2025Efficient . We propose a different representation that enables more efficient computations via our low-rank update technique.
We first rewrite the rank-1 perturbation in the Cholesky-style matrix decomposition form as follows:
@@
Then we can rewrite the design matrix as
@@
$L_{t+1}=L_{t}(I+\alpha_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})\;,\quad\bar{x}_{t+1}=L_{t}^{-1}x_{t+1}\;,$ (13)
@@
$1+\alpha_{t+1}\,\|\bar{x}_{t+1}\|^{2}\;=\;\sqrt{\,1+\|\bar{x}_{t+1}\|^{2}\,}\;.$ (14)
The correctness of the specified recursive definition of $L_{t+1}$ in (13) can be verified by direct substitution. The following proof for the rank-1 updates setting will be useful for illustrating the essence of our approach.
@@
Let $\varepsilon>0$, $L_{0}=\varepsilon^{-1/2}I\in\mathbb{R}^{d\times d}$, and $U_{0}$, $V_{0}$ be empty matrices. Given a sequence of context vectors $\{x_{t}\}_{t\in\mathbb{N}}$ set
$\beta_{t+1}\;=\;\frac{\alpha_{t+1}}{1+\alpha_{t+1}\,\|\bar{x}_{t+1}\|^{2}}.$ (15)
@@
$U_{t+1}\;=\;\begin{bmatrix}U_{t}&\beta_{t+1}\,\bar{x}_{t+1}\end{bmatrix},\qquad V_{t+1}\;=\;\begin{bmatrix}V_{t}&(I-V_{t}\,U_{t}^{\top})\,\bar{x}_{t+1}\end{bmatrix}$. (17)
@@
Using (13), we obtain
$L_{t+1}^{-1}=(I+\alpha_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})^{-1}L_{t}^{-1}\;.$ (18)
After applying the Sherman-Morrison-Woodbury formula
the recursive structure of the inverse update reads:
$L_{t}^{-1}=(I-\beta_{t}\bar{x}_{t}\bar{x}_{t}^{\top})L_{t-1}^{-1}\;.$ (20)
Unraveling this recursion from $t$ back to the initial condition gives
$L_{t+1}^{-1}$ $=(I-\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})L_{t}^{-1}$ (23)
This can be written as $L_{t+1}^{-1}=(I-U_{t+1}V_{t+1}^{\top})L_{0}^{-1}$, where
$U_{t+1}$ $=\begin{bmatrix}U_{t}&\beta_{t+1}\,\bar{x}_{t+1}\end{bmatrix}$, (27)
This completes the proof. ∎

### v2

@@ Abstract
In this paper, we introduce PSI-LinUCB, a scalable variant of LinUCB that enables efficient training, inference, and memory usage by representing the inverse regularized design matrix as a sum of a diagonal matrix and low-rank correction. We derive numerically stable rank-1 and batched updates that maintain the inverse without explicitly forming the matrix. To control memory growth, we employ a projector-splitting integrator for dynamical low-rank approximation, yielding an average per-step update cost and memory usage of $\mathcal{O}(dr)$ for approximation rank $r$. The inference complexity of the proposed algorithm is $\mathcal{O}(dr)$ per action evaluation. Experiments on recommender system datasets demonstrate the effectiveness of our algorithm.
@@
1 Introduction and Problem Setting
Contextual bandits are essential in modern decision making, as they enable online adaptation to dynamic environments and explicitly balance exploration and exploitation auer2002finite. These methods are well studied theoretically auer2002finite, abbasi2011improved, agrawal2013thompson and are widely used in practice, in particular, in the real-world recommender systems li2010contextual, yi2023online. Within this class of methods, LinUCB li2010contextual, abbasi2011improved is one of the most commonly used algorithms. It models the expected reward of an action as a linear function of a $d$-dimensional context vector and follows the Upper Confidence Bound (UCB) principle, implementing optimism in the face of uncertainty auer2002finite. LinUCB admits practical implementations for online recommendation settings li2010contextual and serves as a natural baseline when online updates are required.
@@ 1 Introduction and Problem Setting
For the most part of this paper, we focus on the disjoint LinUCB parametrization li2010contextual, das2024linear, yet the techniques that we develop further naturally generalize to hybrid and shared parametrizations used in this algorithm. Formally, at each time step $t\in\{1,\ldots,T\}$, agent observes a set of arms (actions) $\mathcal{A}_{t}$, where each arm $a\in\mathcal{A}$ is associated with a context vector ${x}_{t,a}\in\mathbb{R}^{d}$. LinUCB assumes a linear reward model,
@@
$a_{t}=\arg\max_{a\in\mathcal{A}_{t}}\left(\hat{{\theta}}_{a}^{\top}{x}_{t,a}+\alpha\sqrt{{x}_{t,a}^{\top}A_{t,a}^{-1}{x}_{t,a}}\right),$ (6)
@@
where $\alpha\in\mathbb{R}$ controls the exploration–exploitation trade-off. In online learning problems $\alpha>0$, but sometimes learning problems with fixed dataset require setting negative values of $\alpha$ (the so-called anti-exploration), see e.g. rezaeifar2022offline, xu2024provably.
LinUCB can also be extended to batch-update settings, where model parameters are updated after multiple interactions (see Appendix˜A in Appendix). At the same time, LinUCB algorithm has several scalability challenges:
@@
  • Matrix inversion: The complexity of computing $A_{t,a}^{-1}$ is $O(d^{3})$, which makes the algorithm computationally expensive if context dimension $d$ is large.
@@
  • Large action space: The need to store and update a separate $d\times d$ matrix for each action leads to increased time and memory requirements as the number of actions grows.
Existing work addresses the scalability limitations of LinUCB in several ways. The simplest acceleration relies on exact rank-one updates of the inverse design matrix using the Sherman–Morrison formula Angioli2025Efficient, which avoids full matrix inversion but still requires maintaining dense $d\times d$ matrices and does not scale to large feature dimensions and large action space. Another line of work focuses on approximating the design matrix $\sum_{s=1}^{t}x_{s,a}x_{s,a}^{\top}$ with a low-rank representation using matrix sketching techniques kuzborskij2019efficient, chen2020efficient, wen2024matrix. While these methods significantly reduce memory and computation costs, they typically process observations sequentially and do not naturally support vectorized or batch updates. An alternative approach applies random feature projections to reduce the dimensionality of the context vectors before learning yu2017cbrap. Such methods enable faster updates but often require a relatively large projected dimension to preserve recommendation quality, making them less efficient than sketching-based low-rank approximations. Finally, diagonal or block-diagonal approximations of the design matrix have been proposed to reduce computational and memory costs yi2023online. We provide a detailed overview of the existing literature in the Appendix, Appendix˜B.
Our contrubutions. In this paper, we propose a scalable variant of LinUCB designed for large-scale contexts settings. Our approach is based on a suitable representation of the inverse regularized design matrix. Specifically, we approximate $A_{t,a}^{-1}$ as the sum of a diagonal matrix and a low-rank correction. This representation avoids explicit matrix inversion, and significantly reduces memory usage by storing only the diagonal terms and low-rank factors. Our primary contributions are as follows:
  • We introduce PSI-LinUCB, a scalable LinUCB variant that maintains a Cholesky-style representation of the inverse regularized design matrix as a sum of diagonal term and dynamically updated low-rank component. Using a projector-splitting method lubich2014projector, our algorithm avoids explicit matrix inversion. Our method naturally supports vectorized and batch updates over multiple context vectors, which is crucial for efficient deployment in modern recommender systems. PSI-LinUCB achieves an average per-interaction update cost of $\mathcal{O}(dr)$ under the proper choice of batch size, with $r$ being the approximation rank. The inference complexity is $\mathcal{O}(dr)$ per action.
  • We empirically show that PSI-LinUCB is robust with respect to the choice of the approximation rank: increasing $r$ leads to a smooth and nearly linear improvement in recommendation quality (measured by hit rate), making the method easy to tune in practice.
  • We evaluate PSI-LinUCB on several large-scale datasets from recommender systems domain, where standard LinUCB fails to run due to scalability constraints. We ensure reduction in memory consumption, and computational time compared to the exact Sherman–Morrison implementation of LinUCB Angioli2025Efficient, while matching the hit rate of the vanilla LinUCB algorithm. We also show that PSI-LinUCB outperforms sketching-based baselines in terms of computational time, while achieving similar or better quality, both on large-scale datasets and in online synthetic environments.
@@ 2 Scalable LinUCB with Low-Rank Updates
The main computational cost of LinUCB lies in updating the parameter estimate $\hat{\theta}_{a}$ for each arm, which requires access to the inverse matrix $A_{t,a}^{-1}$. Existing approaches mitigate this issue either by applying rank-1 Sherman–Morrison updates Angioli2025Efficient or by approximating the covariance matrix
@@
using low-rank sketching techniques such as Frequent Directions kuzborskij2019efficient and CBSCFD chen2020efficient. In contrast to these works, we employ an alternative approach, which relies on the dynamic representation of the inverse matrix $A_{t,a}^{-1}$. Towards this aim, we use the Cholesky-style symmetric factorization $A_{t,a}=L_{t,a}L_{t,a}^{\top}$ and note that in this case
$A_{t,a}^{-1}=L_{t,a}^{-\top}L_{t,a}^{-1}\;.$ (8)
The motivation for working with the inverse design matrix is that in LinUCB the design matrix $A_{t,a}$ is never used directly: both the parameter estimate
$\hat{\theta}_{t,a}=A_{t,a}^{-1}b_{t,a}$ (9)
and the exploration bonus
$\sqrt{x_{t,a}^{\top}A_{t,a}^{-1}x_{t,a}}$ (10)
involve only matrix–vector products with $A_{t,a}^{-1}$. Therefore, maintaining or approximating $A_{t,a}$ itself is algorithmically unnecessary.
The Cholesky-style factorization (8) makes this observation particularly convenient. Indeed, it allows both quantities above to be computed via matrix-vector operations involving $L_{t,a}^{-1}$ only. This representation enables us to work directly with the inverse operator without explicitly forming or storing $A_{t,a}^{-1}$.
The representation (8) is a key formula for our further analysis. Importantly, we do not explicitly compute the inverse factors $L_{t,a}^{-1}$. Instead, they are updated dynamically from their previous values. Since the update rules are identical across arms, we omit the index $a$ in the remainder of this section and write $A_{t}$ and $L_{t}$ (respectively, $L_{t}^{-1}$) instead of $A_{t,a}$ and $L_{t,a}$.
In the next parts of this section, we present our methodology in different setups. First, in Section˜2.1 we consider the case of rank-1 updates and write the dynamics of updates of the inverse root matrix $L_{t}^{-1}$. Then in Section˜2.2 we generalize our expressions for the batch update case. Both representations rely on a recursive definition in the form of matrix decomposition $U_{t}V_{t}^{\top}$. In Section˜3, we show that this decomposition can be dynamically updated in a low-rank format using the Projector-Splitting Integrator (PSI) method of Lubich and Oseledets lubich2014projector. Such updating mechanism prevents an uncontrollable growth of factor matrices $U_{t}$ and $V_{t}$ over time and enables “on-the-fly” adaptation to the stream of contextual data. This approach has been previously considered in the context of modifying the PureSVD model in recommender systems Olaleke2021Dynamic.
@@ 2.1 Rank-1 updates
We first consider the rank-1 update setting, where context vectors arrive sequentially and model parameters are updated after each interaction. In this case, the design matrix update at time $t+1$ takes the form
$A_{t+1}=A_{t}+x_{t+1}x_{t+1}^{\top}.$ (11)
While inverse updates can be obtained via the Sherman–Morrison formula Angioli2025Efficient, we instead derive an alternative representation that enables efficient low-rank updates within our proposed framework.
We begin by expressing the rank-1 update in a Cholesky-style factorized form. Using $A_{t}=L_{t}L_{t}^{\top}$, the update at time $t+1$ can be written as
@@
Hence, we can rewrite the design matrix as
@@
$L_{t+1}=L_{t}(I+\alpha_{t+1}\tilde{x}_{t+1}\tilde{x}_{t+1}^{\top})\;,\;\tilde{x}_{t+1}=L_{t}^{-1}x_{t+1}\;,$ (14)
@@
$1+\alpha_{t+1}\,\|\tilde{x}_{t+1}\|^{2}\;=\;\sqrt{\,1+\|\tilde{x}_{t+1}\|^{2}\,}\;.$ (15)
The correctness of the specified recursive definition of $L_{t+1}$ in (14) can be verified by direct substitution. The following theorem for the rank-1 updates setting will be useful for illustrating the essence of our approach.
@@
Let $\varepsilon>0$, $L_{0}^{-1}=\varepsilon^{-1/2}I\in\mathbb{R}^{d\times d}$, and $U_{0}$, $V_{0}$ be empty matrices. Given a sequence of context vectors $\{x_{t}\}_{t\in\mathbb{N}}$ set
$\beta_{t+1}\;=\;\frac{\alpha_{t+1}}{1+\alpha_{t+1}\,\|\tilde{x}_{t+1}\|^{2}}.$ (16)
@@
$U_{t+1}=\begin{bmatrix}U_{t}&\beta_{t+1}\,\tilde{x}_{t+1}\end{bmatrix},V_{t+1}=\begin{bmatrix}V_{t}&(I-V_{t}\,U_{t}^{\top})\,\tilde{x}_{t+1}\end{bmatrix}$. (18)
@@
The proof is provided in the Appendix, Section˜D.1